OpenStudy (anonymous):

How many solutions does the system of equations have? y - 7x = -14 and 7y - 49x = -2 one two infinite no solution

OpenStudy (luigi0210):

Let's get them into slope-intercept form.

OpenStudy (anonymous):

okay I know how to do that

OpenStudy (luigi0210):

Okay, so put each in \[\LARGE y=mx+b\] and show me what you get.

OpenStudy (anonymous):

the slope intercept form of 7y - 49x = -2 is y = 7x - 2/7

OpenStudy (luigi0210):

Yea, now what about the first equation?

OpenStudy (anonymous):

y = 7x - 14

OpenStudy (luigi0210):

Correct, so now we have: \[\LARGE y=7x-14\]and \[\LARGE y=7x-\frac{2}{7}\] Notice anything about them> Anything they have in common?

OpenStudy (anonymous):

the slope is the same

OpenStudy (anonymous):

If they have the same slope but different y-intercepts, there is no solution correct?

OpenStudy (luigi0210):

Exactly! And now that we have this information, we are left with deciding whether they have no solution or infinitely many. And yea, that's right :) They are parallel, so they never cross. Nice job :)

OpenStudy (anonymous):

Thank you! (: I have a hard time remembering these rules, is there a key or chart of some sort?

OpenStudy (luigi0210):

Hmm, I don't have much.. just this:

OpenStudy (luigi0210):

But the key would be looking at the equations.

OpenStudy (anonymous):

Okay! thanks so much! (:

OpenStudy (luigi0210):

Anytime, and welcome to the site ^_^